# Mitigate the Gap: Investigating Approaches for Improving Cross-Modal Alignment in CLIP

### v1

@@ 2 Background, Notations and Concepts
$\mathcal{L}_{v\rightarrow l}=\frac{1}{N}\sum_{i=1}^{N}\log\frac{\exp[(\vec{e}_{\mathrm{v}}^{\,i}\cdot\vec{e}_{\mathrm{t}}^{\,i})\mathbin{/}\tau]}{\sum_{j=1}^{N}\exp[(\vec{e}_{\mathrm{v}}^{\,i}\cdot\vec{e}_{\mathrm{t}}^{\,j})\mathbin{/}\tau]},\quad\mathcal{L}_{l\rightarrow v}=\frac{1}{N}\sum_{j=1}^{N}\log\frac{\exp[(\vec{e}_{\mathrm{v}}^{\,j}\cdot\vec{e}_{\mathrm{t}}^{\,j})\mathbin{/}\tau]}{\sum_{i=1}^{N}\exp[(\vec{e}_{\mathrm{v}}^{\,i}\cdot\vec{e}_{\mathrm{t}}^{\,j})\mathbin{/}\tau]}$ (1)
, respectively, where $\tau$ is the learnable temperature parameter. The overall CLIP loss is then:
@@
In practice, a symmetric cross-entropy loss is employed using the vision and language logits. The label $y\in\mathbb{R}$, which is the index of the image-text pair in the batch, represents the correspondence of the paired samples. For a batch of image–text pairs, $Y\in\mathbb{R}^{b}$ denotes the set of labels. Denoting $\langle A,B\rangle=A\cdot B^{\intercal}$, the visual and textual logits, $\hat{y}_{\mathrm{v}}\in\mathbb{R}^{b\times b}$ and $\hat{y}_{\mathrm{t}}\in\mathbb{R}^{b\times b}$, are then calculated as:
$\hat{y}_{\mathrm{v}}=\exp(\tau)\cdot\langle E_{\mathrm{v}},E_{\mathrm{t}}\rangle\quad,\quad\hat{y}_{\mathrm{t}}=\hat{y}_{\mathrm{v}}^{\intercal},$ (3)
@@ 3.2 Intra-Modality Separation
$\mathcal{V}=\langle E_{\mathrm{v}}\cdot E_{\mathrm{v}}\rangle,\quad\mathcal{T}=\langle E_{\mathrm{t}}\cdot E_{\mathrm{t}}\rangle,$ (7)
@@
$\mathcal{S}=\frac{\langle E_{\mathrm{s}}\cdot E_{\mathrm{s}}\rangle}{\left\lVert E_{\mathrm{s}}\right\rVert^{2}},\quad\quad\mathcal{D}=1-\mathcal{S},$ (8)
@@
where $\odot$ denotes the element-wise product, in which $\mathcal{V_{D}}\in\mathbb{R}^{b\times b}$ and $\mathcal{T_{D}}\in\mathbb{R}^{b\times b}$. Afterwards, we calculate $\mathcal{M}=\langle E_{\mathrm{v}},E_{\mathrm{t}}\rangle$ and mask the non-diagonal values by $\textrm{diag}(\mathcal{M})=\mathbb{I}\odot\mathcal{M}$. We then obtain the vision-vision logits $\hat{y}_{\mathrm{vsep}}\in\mathbb{R}^{b\times b}$ as well as text-text logits $\hat{y}_{\mathrm{tsep}}\in\mathbb{R}^{b\times b}$ by:
@@ 4.1 Training Dataset and Setup
We used the Conceptual Caption 12M (CC12M) dataset, which has been similarly used in previous work (Goel et al., 2022; Mu et al., 2022; Li et al., 2021; Changpinyo et al., 2021), for pre-training the models. We set our setup quite similar to that of the original CLIP with the ViT-B-16 backend, in order to ensure a fair comparison of CLIP with SharedCLIP and AlignCLIP. The pre-trained semantic encoder utilized in AlignCLIP for re-scaling text–text and image–image cosine similarities is the SBERT all-mpnet-base-v2 model. In order to fairly compare the effectiveness of each model, we trained all of them from scratch using the CC12M dataset and the OpenCLIP implementation (Cherti et al., 2023; Ilharco et al., 2021). Each model was trained for 6 days using an NVIDIA H100 GPU with batch size 512 for 30 epochs using AdamW optimization with a starting learning rate of $1\times 10^{-3}$, cosine scheduler, 10,000 warmup steps, a weight decay of $0.1$, and an initial temperature value of 0.07. The output embedding dimensionality for both the vision and language modalities is set to 768. We used the checkpoint from the last epoch in our evaluations of downstream experiments. In AlignCLIP, we set $\alpha=1$ and $\beta=\frac{1}{2}$. In terms of ablation study, we also trained IMSep loss without sharing the parameters between vision and language encoders. In Section A.1, more details about the training setup is provided. Furthermore, in Section A.3, we repeat our experiment when using output embedding dimension of 512 and show the results.

### v2

@@ 2 Background, Notations and Concepts
$\mathcal{L}_{v\rightarrow l}=\frac{1}{N}\sum_{i=1}^{N}\log\frac{\exp[(\vec{e}_{\mathrm{v}}^{\,i}\cdot\vec{e}_{\mathrm{t}}^{\,i})\mathbin{/}\tau]}{\sum_{j=1}^{N}\exp[(\vec{e}_{\mathrm{v}}^{\,i}\cdot\vec{e}_{\mathrm{t}}^{\,j})\mathbin{/}\tau]},\quad\mathcal{L}_{l\rightarrow v}=\frac{1}{N}\sum_{j=1}^{N}\log\frac{\exp[(\vec{e}_{\mathrm{v}}^{\,j}\cdot\vec{e}_{\mathrm{t}}^{\,j})\mathbin{/}\tau]}{\sum_{i=1}^{N}\exp[(\vec{e}_{\mathrm{v}}^{\,i}\cdot\vec{e}_{\mathrm{t}}^{\,j})\mathbin{/}\tau]},$ (1)
respectively, where $\tau$ is the learnable temperature parameter. The overall CLIP loss is then:
@@
In practice, a symmetric cross-entropy loss is employed using the vision and language logits. The label $y\in\mathbb{R}$, which is the index of the image-text pair in the batch, represents the correspondence of the paired samples. For a batch of image–text pairs, $Y\in\mathbb{R}^{b}$ denotes the set of labels. The visual and textual logits, $\hat{y}_{\mathrm{v}}\in\mathbb{R}^{b\times b}$ and $\hat{y}_{\mathrm{t}}\in\mathbb{R}^{b\times b}$, are then calculated as:
$\hat{y}_{\mathrm{v}}=\exp(\tau)E_{\mathrm{v}}E_{\mathrm{t}}^{\intercal}\quad,\quad\hat{y}_{\mathrm{t}}=\hat{y}_{\mathrm{v}}^{\intercal},$ (3)
@@ 3.2 Intra-Modality Separation
$\mathcal{V}=E_{\mathrm{v}}E_{\mathrm{v}}^{\intercal},\quad\mathcal{T}=E_{\mathrm{t}}E_{\mathrm{t}}^{\intercal},$ (7)
@@
$\mathcal{S}=\frac{E_{\mathrm{s}}E_{\mathrm{s}}^{\intercal}}{\left\lVert E_{\mathrm{s}}\right\rVert^{2}},\quad\quad\mathcal{D}=1-\mathcal{S},$ (8)
@@
where $\odot$ denotes the element-wise product, in which $\mathcal{V_{D}}\in\mathbb{R}^{b\times b}$ and $\mathcal{T_{D}}\in\mathbb{R}^{b\times b}$. Afterwards, we calculate $\mathcal{M}=E_{\mathrm{v}}E_{\mathrm{t}}^{\intercal}$ and mask the non-diagonal values by $\textrm{diag}(\mathcal{M})=\mathbb{I}\odot\mathcal{M}$. We then obtain the vision–vision logits $\hat{y}_{\mathrm{vsep}}\in\mathbb{R}^{b\times b}$ as well as text–text logits $\hat{y}_{\mathrm{tsep}}\in\mathbb{R}^{b\times b}$ by:
@@ 4.1 Training Dataset and Setup
We used the Conceptual Caption 12M (CC12M) dataset, which has been similarly used in previous work (Goel et al., 2022; Mu et al., 2022; Li et al., 2022; Changpinyo et al., 2021), for pre-training the models. We set our setup quite similar to that of the original CLIP with the ViT-B-16 backend, in order to ensure a fair comparison of CLIP with SharedCLIP and AlignCLIP. The pre-trained semantic encoder utilized in AlignCLIP for re-scaling text–text and image–image cosine similarities is the SBERT all-mpnet-base-v2 model. In order to fairly compare the effectiveness of each model, we trained all of them from scratch using the CC12M dataset and the OpenCLIP implementation (Cherti et al., 2023; Ilharco et al., 2021). Each model was trained for 6 days using an NVIDIA H100 GPU with batch size 512 for 30 epochs using AdamW optimization with a starting learning rate of $1\times 10^{-3}$, cosine scheduler, 10,000 warmup steps, a weight decay of $0.1$, and an initial temperature value of 0.07. The output embedding dimensionality for both the vision and language modalities is set to 768. We used the checkpoint from the last epoch in our evaluations of downstream experiments. In AlignCLIP, we set $\alpha=1$ and $\beta=\frac{1}{2}$. In terms of ablation study, we also trained IMSep loss without sharing the parameters between vision and language encoders. In Section A.1, more details about the training setup is provided. Furthermore, in Section A.3, we repeat our experiment when using output embedding dimension of 512 and show the results.